OpenStudy (anonymous):

Segment AB has endpoints A(3,6) and B(9,-2) and midpoint M. Find the coordinates of M and Find AB.

OpenStudy (alexandervonhumboldt2):

use midpoint formula: \[(\frac{ x_1+x_2 }{ 2 }. \frac{ y_1+y_2 }{ 2 })\]

OpenStudy (alexandervonhumboldt2):

substitute your points and solve

OpenStudy (anonymous):

Thank you @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

for AB use distance formula: \[\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\]